OpenStudy (anonymous):

Solve using the quadratic forumula. 3x^2 + x - 4 = 0

OpenStudy (anonymous):

Do you know the quadratic formula? \[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

Then it's just a matter of plugging in the values and simplifying.

OpenStudy (anonymous):

3x^2+4x-3x-4=0 (x-1)(3x+4)=0

OpenStudy (anonymous):

x=1,-4/3

OpenStudy (anonymous):

I understand how to substitute everything in, but from there I am unsure how to simplify -1 ± 12−4*3*4−−−−−−−√

OpenStudy (anonymous):

-1 +/- sqrt -48

OpenStudy (anonymous):

a=3,b=1, c=-4

OpenStudy (anonymous):

Work on the square root part first. \[\sqrt{1^{2}-4*3*-4}\]

OpenStudy (anonymous):

You missed a couple things in the sqrt. \[1^{2}=1\] 4*3*-4=-48 1-(-48)

OpenStudy (anonymous):

Get that? So it's \[\sqrt{1+48}=\sqrt{49}=7\] So, \[(1 \pm 7)/6\]

OpenStudy (anonymous):

Thank you very much! That makes much more sense now. Once again, thanks!

OpenStudy (anonymous):

np.